### 31 Jan 2024 v1

@@ 4.4 Weak activity period - type-IV burst
The post-flare weak activity phase is found to last for 1.4 h. The Stokes V dynamic spectrum (Fig. 1a) shows that the post-flare emission shows a bright patch confined to frequencies 700 MHz, between 2021-12-03 23:27 and 2021-12-04 02:25 UT. To highlight this better a 200 s averaged VISAD spectrum was made (see, Fig. 6a). The imaging-based spectrum from the weak activity period (Fig. 3b) also shows a characteristic high polarisation bright emission feature above 700 - 730 MHz, where the emission polarisation jumps above 55% and rises to 80% as frequency increases. As per the solar radio burst classification scheme, the aforementioned features suggest that the weak activity phase corresponds to a type-IV radio burst confined to frequencies above 700 MHz (Gopalswamy, 2011; Hillaris et al., 2016; Salas-Matamoros & Klein, 2020). For comparison, Fig. 6b shows a solar dynamic spectrum during a strong CME taken by Wind/WAVES instrument. The solar data is taken from the publicly available type-IV catalogue compiled by Coordinated Data Analysis Workshops (CDAW) data center 333https://cdaw.gsfc.nasa.gov/CME_list/radio/type4. The solar event was so chosen to match the dynamic spectral emission profile of the AD Leo. The morphological analogues of the type-III and type-IV bursts in the solar template spectrum can be identified in the AD Leo DS.
@@
Type-IV radio bursts are associated with strong flares and CMEs in the Sun, like the much sought-after type-II bursts (Robinson, 1986; Gopalswamy, 2011; Hillaris et al., 2016; Kumari et al., 2021). While a type-II burst is associated with the coronal shock generated by a fast CME, a type-IV burst is generated by flare-accelerated electrons trapped either in the post-flare loops or in the moving plasmoids associated with the CME (e.g., Wild, 1970; Smerd, 1970; Gopalswamy et al., 2005; White, 2007; Gopalswamy, 2011). Type-IVs are generally associated with relatively longer and stronger solar X-ray flares than type-IIs (Cane & Reames, 1988). Also, type-IV bursts are often linked to strong CMEs which expand out to heliocentric angular widths 60, with mean speeds 900 km s (Robinson, 1986; Gopalswamy, 2011; Hillaris et al., 2016). Recently, Zic et al. (2020) reported the first type-IV burst detection in a star other than the Sun, namely Proxima Centauri which is an old (5 Gyr) M dwarf. Given its age ( 250 Myr) and rotation period (P = 2.2 days), AD Leo belongs to a fast rotating (period 5 days), young (age 1 Gyr) and more active stellar population (‘C’ branch) compared to the the Sun (age = 4.6 Gyr; P = 25 days), which belongs to the less active ‘I’ branch stars (Barnes, 2003; Brown, 2014). Barnes (2003) demonstrated that the stars undergo an evolutionary migration from ‘C’ to ‘I’ branch as they spin down with age due to mass and angular momentum loss. This migration involves a fundamental change in the way the internal dynamo couples with the external active atmospheric layers and the stellar wind (Brown, 2014; Garraffo et al., 2018). So, though Proxima Centauri is an M dwarf, it is an ‘I’ branch star, but AD Leo is not. Here, we report the first detection of a type-IV burst on a young (250 Myr), active ‘C’ branch M dwarf.
@@
We do not find evidence for a type-II burst, making it still elusive in an M dwarf despite several day-long monitoring campaigns on various stars, including AD Leo (e.g., Osten et al., 2008; Villadsen & Hallinan, 2019). Though rare, solar type-II bursts have also been found associated with coronal shocks produced by failed eruptions, unrelated to CMEs (e.g., Magdalenić et al., 2012; Morosan et al., 2023). AD Leo is an active star with an order of magnitude stronger and predominantly axi-symmetric magnetic field (Moschou et al., 2019). Theoretical models have argued that the strong magnetic field structures and the associated high Alfvén speed barriers in the stellar corona could either block even the energetic flares from causing an eruption/CME (failed eruption scenario), or resist the formation of a strong shock required to initiate a type-II emission during an eruption/CME (Odert et al., 2020; Alvarado-Gómez et al., 2022). This means that, unlike the case of the sun, AD Leo could host more failed eruptions during strong flares, and even if a CME/eruption occurs, type-II bursts need not be triggered for most events. These scenarios can explain why have there been no report of a type-II burst on AD Leo despite several monitoring campaigns (e.g., Osten & Bastian, 2006; Osten et al., 2008; Villadsen & Hallinan, 2019). In any case, since post-flare loop structures will trap flare-accelerated electrons, strong type-IV bursts could trace CMEs better than type-IIs, in M dwarfs.
@@
Given the high X-mode polarisation level and high at 700 MHz, and based on our inferences for the active region in the previous section, the type-IV burst emission likely originates from an ECME mechanism in the Z/W mode coalescence or a harmonic emission pathways. An ECME mechanism is usually triggered by an unstable loss-cone distribution of energetic electrons formed within a magnetic bottle like configuration, readily provided by the post-flare loops (e.g., Carley et al., 2019). In the case of the sun, type-IV bursts are generally associated with plasma emission mechanism in the metric waveband (Weiss, 1963; Salas-Matamoros & Klein, 2020). However, during some strong flares, associated type-IV bursts have demonstrated ms to s scale spiky emission features in their dynamic spectra with high degree of circular polarisation and , often modelled as generated by an ECME mechanism (e.g., Carley et al., 2019; Liu et al., 2018; Morosan et al., 2019; Cliver et al., 2022). Certain high time resolution studies of AD Leo in 1-3 GHz observations have reported ms scale spiky emission features (e.g., Osten et al., 2008; Zhang et al., 2023). The time resolution of uGMRT data restricts us from exploring sub-second scale fine structures in the dynamic spectrum. The ECME signatures we detect in the type-IV emission above 700 MHz could probably be formed by numerous sub-second scale fine structures. Meanwhile, at frequencies below 700 MHz the emission polarisation is well below 30%. The emission mechanism here is likely a H-branch plasma emission which is known to produce high emission at low polarisation levels (e.g., Willes & Melrose, 1997; Saint-Hilaire et al., 2013; Rahman et al., 2019).
@@
Since the type-IV emission is seen to be mostly confined within 700 MHz (Fig. 6), using Model 2 and assuming that the ECME is formed at either the second or the third harmonic of , the height of the post flare loop can be inferred. Consider Fig. 5. It can be seen from Fig. 5b that within 1.8 - 2.25 R, the overlap region of cyan and gray bands satisfy the condition of 0.3¡/¡1.5 required for Z-mode/harmonic X-mode mechanism to operate efficiently and produce a second harmonic ECME observable in . Since 700 MHz is roughly the centre of cyan band, the post-flare loop height could be 1.8 - 2.0 R. Similarly, if the emission is at the third harmonic (= 3), we should use Fig. 4(a-b) to obtain emission height. Here instead of the cyan band, we consider the overlap between the green and the gray bands. The emission height range turns out to be 2.0 -2.5 R, implying a post-flare loop height of 2 - 2.25 R. This implies that the star can have loops rising to about 1R above the photosphere.
@@ 4.5 Quasi-steady period
The quiescent Stokes I flux level varies within 1 - 1.75 mJy (see Figs. 1c & 2) implying values of the order of 10 K. However, the polarisation levels of 20 - 30% inferred from the light curve suggests a H-band coherent plasma emission mechanism. The observed emission is likely a superposition of various spatially separated coherent plasma emission regions or a mixture of gyrosynchroton and plasma emission regions (Bastian et al., 1990; Bastian, 1994).
@@ 5 Conclusions
We present a two hour-long metric band flaring activity in a young active M dwarf, AD Leo. A hydrostatic equilibrium electron density profile was assumed with a basal density ranging within 10 - 10 cm. Based on the existing knowledge on the stellar magnetic field structure, different radial field profiles (B(r)) were considered: a solar-like model and a simple multipole expansion model. Contrasting the radio data with these models we conclude the following:
Active region B(r) is more consistent with a multipole expansion model than a solar-like model, given the possible range of B(r) and basal coronal density.
@@
The period of strong flaring was succeeded by a post-flare activity period. We report a solar like type-IV radio burst in AD Leo, in the post-flare phase characterised by a distinctive dynamic spectral morphology and circular polarisation levels ranging within 60 - 100%. Previously, a type-IV emission was reported in Proxima Centauri by Zic et al. (2020). This is the first solar-like type-IV detection on a young active M dwarf belonging to a different age-activity population (‘C’ branch) relative to the Sun (belonging to ‘I’ branch) (Barnes, 2003). The emission characteristics favour an ECME mechanism at the second or third harmonic of the local gyrofrequency. The detection of the type-IV burst could signify a strong CME, since solar type-IVs are known to be associated with fast and wide CMEs. Nevertheless, an hour-long type-IV burst signifies powerful accelerated particle activity in the post-flare loops. Based on the frequency extent of the type-IV burst in the VISAD dynamic spectrum we infer that the post-flare loops could rise up to 1 R scales above photosphere into the inter-planetary space. The quasi-steady emission that followed the type-IV activity is likely powered by a mixture of several coherent plasma emission regions.

### 04 Mar 2024 v2

@@ 4.4 Weak activity period and the type-IV burst
The post-flare weak activity phase is found to last for 1.4 h. The Stokes V dynamic spectrum (Fig. 1a) shows that the post-flare emission shows a bright patch confined to frequencies 700 MHz, between 2021-12-03 23:27 and 2021-12-04 02:25 UT. To highlight this better, a 200 s averaged VISAD spectrum was made (see Fig. 6a). The imaging-based spectrum from the weak activity period (Fig. 3b) also shows a characteristic high-polarisation bright-emission feature above 700 - 730 MHz, where the emission polarisation jumps above 55% and rises to 80% as frequency increases. As per the solar radio burst classification scheme, the aforementioned features suggest that the weak activity phase corresponds to a type-IV radio burst confined to frequencies above 700 MHz (Gopalswamy, 2011; Hillaris et al., 2016; Salas-Matamoros & Klein, 2020). For comparison, Fig. 6b shows a solar dynamic spectrum during a strong CME taken by the Wind/WAVES instrument. The solar data are taken from the publicly available type-IV catalogue compiled by Coordinated Data Analysis Workshops (CDAW) data centre 333https://cdaw.gsfc.nasa.gov/CME_list/radio/type4. The solar event was so chosen to match the dynamic spectral emission profile of the AD Leo. The morphological analogues of the type-III and type-IV bursts in the solar template spectrum can be identified in the AD Leo DS.
@@
Type-IV radio bursts are associated with strong flares and CMEs in the Sun, as the sought-after type-II bursts (Robinson, 1986; Gopalswamy, 2011; Hillaris et al., 2016; Kumari et al., 2021). While a type-II burst is associated with the coronal shock generated by a fast CME, a type-IV burst is generated by flare-accelerated electrons trapped either in the post-flare loops or in the moving plasmoids associated with the CME (e.g. Wild, 1970; Smerd, 1970; Gopalswamy et al., 2005; White, 2007; Gopalswamy, 2011). Type-IVs are generally associated with relatively longer and stronger solar X-ray flares than type-IIs (Cane & Reames, 1988). Also, type-IV bursts are often linked to strong CMEs, which expand out to heliocentric angular widths 60, with mean speeds 900 km s (Robinson, 1986; Gopalswamy, 2011; Hillaris et al., 2016). Recently, Zic et al. (2020) reported the first type-IV burst detection in a star other than the Sun, namely Proxima Centauri, which is an old (5 Gyr) M dwarf. Given its age ( 250 Myr) and rotation period (P = 2.2 days), AD Leo belongs to a fast rotating (period 5 days), young (age 1 Gyr) and more active stellar population (‘C’ branch) compared to the the Sun (age = 4.6 Gyr; P = 25 days), which belongs to the less active ‘I’ branch stars (Barnes, 2003; Brown, 2014). Barnes (2003) demonstrated that the stars undergo an evolutionary migration from ‘C’ to ‘I’ branch as they spin down with age due to mass and angular momentum loss. This migration involves a fundamental change in the way the internal dynamo couples with the external active atmospheric layers and the stellar wind (Brown, 2014; Garraffo et al., 2018). So, though Proxima Centauri is an M dwarf, it is an ‘I’ branch star, but AD Leo is not. Here, we report the first detection of a type-IV burst on a young (250 Myr), active ‘C’ branch M dwarf.
@@
We do not find evidence of a type-II burst, making it still elusive in an M dwarf despite monitoring campaigns spanning several days on various stars, including AD Leo (e.g. Osten et al., 2008; Villadsen & Hallinan, 2019). Though rare, solar type-II bursts have also been found to be associated with coronal shocks produced by failed eruptions, unrelated to CMEs (e.g. Magdalenić et al., 2012; Morosan et al., 2023). AD Leo is an active star with a magnetic field that is an order of magnitude stronger and predominantly axi-symmetric (Moschou et al., 2019). Theoretical models have argued that the strong magnetic field structures and the associated high Alfvén speed barriers in the stellar corona could either block even the energetic flares from causing a CME (failed eruption scenario), or resist the formation of a strong shock required to initiate a type-II emission during an eruption/CME (Odert et al., 2020; Alvarado-Gómez et al., 2022). This means that, unlike the case of the Sun, AD Leo could host more failed eruptions during strong flares, and even if a CME or eruption occurs, type-II bursts need not be triggered for most events. These scenarios can explain why there have been no reports of a type-II burst on AD Leo despite several monitoring campaigns (e.g. Osten & Bastian, 2006; Osten et al., 2008; Villadsen & Hallinan, 2019). In any case, since post-flare loop structures will trap flare-accelerated electrons, strong type-IV bursts could trace CMEs better than type-IIs in M dwarfs.
@@
Given the high X-mode polarisation level and high at 700 MHz, and based on our inferences for the active region in the previous section, the type-IV burst emission likely originates from an ECME mechanism via either a Z/W-mode coalescence or a harmonic emission pathway. An ECME mechanism is usually triggered by an unstable loss-cone distribution of energetic electrons formed within a magnetic bottle-like configuration, readily provided by the post-flare loops (e.g. Carley et al., 2019). In the case of the Sun, type-IV bursts are generally associated with plasma emission mechanism in the metric waveband (Weiss, 1963; Salas-Matamoros & Klein, 2020). However, during some strong flares, associated type-IV bursts have demonstrated millisecond to second scale spiky emission features in their dynamic spectra with high degrees of circular polarisation and , often modelled as generated by an ECME mechanism (e.g. Carley et al., 2019; Liu et al., 2018; Morosan et al., 2019; Cliver et al., 2022). Certain high time-resolution studies of AD Leo in 1-3 GHz observations have reported ms scale spiky emission features (e.g. Osten et al., 2008; Zhang et al., 2023). The time resolution of uGMRT data restricts us from exploring sub-second-scale fine structures in the dynamic spectrum. The ECME signatures we detect in the type-IV emission above 700 MHz could probably be formed by numerous sub-second-scale fine structures. Meanwhile, at frequencies below 700 MHz the emission polarisation is well below 30%. The emission mechanism here is likely a H-branch plasma emission, which is known to produce high emission at low polarisation levels (e.g. Willes & Melrose, 1997; Saint-Hilaire et al., 2013; Rahman et al., 2019).
@@
Since the type-IV emission is seen to be mostly confined within 700 MHz (Fig. 6), using Model 2 and assuming that the ECME is formed at either the second or the third harmonic of , the height of the post flare loop can be inferred. It can be seen from Fig. 5b that within 1.8 - 2.25 R, the overlap region of the cyan and grey bands satisfy the condition of 0.3¡/¡1.5 required for Z-mode or harmonic X-mode mechanism to operate efficiently and produce a second harmonic ECME observable in . Since 700 MHz is roughly the centre of the cyan band, the post-flare loop height could be 1.8 - 2.0 R. Similarly, if the emission is at the third harmonic (= 3), we should use Fig. 4(a-b) to obtain emission height. Here, instead of the cyan band, we considered the overlap between the green and the grey bands. The emission height range turns out to be 2.0 -2.5 R, implying a post-flare loop height of 2 - 2.25 R. This implies that the star can have loops rising to about 1R above the photosphere.
@@ 4.5 Quasi-steady period
The quiescent Stokes I flux level varies within 1 - 1.75 mJy (see Figs. 1c & 2), implying values of the order of 10 K. However, the polarisation levels of 20 - 30% inferred from the light curve suggests a H-band coherent plasma emission mechanism. The observed emission is likely a superposition of various spatially separated coherent plasma emission regions or a mixture of gyrosynchroton and plasma emission regions (Bastian et al., 1990; Bastian, 1994).
@@ 5 Conclusions
We present a two-hour-long metric band flaring activity in a young active M dwarf, AD Leo. A hydrostatic equilibrium electron density profile was assumed with a basal density ranging within 10 - 10 cm. Based on the existing knowledge on the stellar magnetic field structure, different radial field profiles (B(r)) were considered: a solar-like model and a simple multipole expansion model. Contrasting the radio data with these models, we conclude the following:
The active region B(r) is more consistent with a multipole expansion model than a solar-like model, given the possible range of B(r) and basal coronal density.
@@
The period of strong flaring was succeeded by a post-flare activity period. We report a solar-like type-IV radio burst in AD Leo, in the post-flare phase, characterised by a distinctive dynamic spectral morphology and circular polarisation levels ranging from 60% - 100%. Previously, a type-IV emission was reported in Proxima Centauri by Zic et al. (2020). This is the first solar-like type-IV detection on a young active M dwarf belonging to a different age-activity population (‘C’ branch) relative to the Sun (belonging to ‘I’ branch) (Barnes, 2003). The emission characteristics favour an ECME mechanism at the second or third harmonic of the local gyrofrequency. The detection of the type-IV burst could signify a strong CME, since solar type-IVs are known to be associated with fast and wide CMEs. Nevertheless, an hour-long type-IV burst signifies powerful accelerated particle activity in the post-flare loops. Based on the frequency extent of the type-IV burst in the VISAD dynamic spectrum, we infer that the post-flare loops could rise up to 1 R scales above photosphere into the inter-planetary space. The quasi-steady emission that followed the type-IV activity is likely powered by a mixture of several coherent plasma emission regions.